OpenStudy (anonymous):

Solve the system y= -1/3x+2 and x+3y=3 Answers: A: 0,1 B: 1,0 C: 3, 1/3 D: 3/2, 0 E: no solution I got this problem wrong and I have no idea what I did wrong. I even forgot what my answer was. Please help

pooja195 (pooja195):

\[\huge~\rm~ y= -1/3+2 ~~~~~~and~~~~~~~ x+3y=3\] \[\huge~\rm~x+3( -1/3+2 )=3\]

pooja195 (pooja195):

I used the substition method

pooja195 (pooja195):

3(−1/3+2) distribute the 3

OpenStudy (anonymous):

So doing 3 times -1/3 then 3 times 2?

pooja195 (pooja195):

yes

OpenStudy (anonymous):

I got -1 and 6